Steam enters a turbine with an enthalpy of 1292 BTU/lbm. The transferred heat is 13 BTU/lbm. What is the work in BTU/min and in hp for a flow of 2lbm/sec?


1
Expert's answer
2021-09-24T09:25:11-0400

First, we know that:


"flow=2 \frac{lb}{s}
\\ h_1=1292\frac{BTU}{lb}; h_2=1098\frac{BTU}{lb}
\\ q=13\frac{BTU}{lb}"


Then we use the energy balance:


"h_1-q=h_2+W
\\ \implies W=h_1-(h_2+q)
\\ W=1292\frac{BTU}{lb}-(1098\frac{BTU}{lb}+13\frac{BTU}{lb}) 
\\ \text{ }
\\ W= 181 \frac{BTU}{lb} \times \frac{2\,lb}{s}\times \frac{60\,s}{1\,min}
\\ \text{ }
\\ W=21720 \frac{BTU}{min} \times \frac{1\,hp}{42.4\,BTU/min}=512.264\,hp"


In conclusion, we find that the work in BTU/min and in hp is 21720 BTU/min and 512.264 hp, respectively.


Reference:

  • Sta. Maria, Hipolito B., Thermodynamics 1, National Bookstore, Inc., Philippines, 2005.